OpenStudy (anonymous):

HELP PLEASE what is the length of the conjugate axis(x-2)^2/36-(y+1)^2/64=1

OpenStudy (anonymous):

)A chord passing through the focus of the parabola x^2=16y has one end at ... The axis of symmetry is x=0, the y axis, because y has the same value for x .... a^2=64 c=6 (distance from origin to foci on the major axis) c^2=36 ... x^2/28+y^2/64=1 .... We can see that the transverse axis is 5 units long.

OpenStudy (anonymous):

( j = 1,..., 6 ) a set of functions defined over the surface of a slender body (see (2-75)) m (x) added mass per unit length of a slender body n. ... The vertical axis is the y axis in 2-D problems, the z axis in 3-D problems. ...... automatically arranges velocities in the anticipated order: 1) Forward speed 2) Transverse perturbation 3)